OpenStudy (anonymous):

The “Limited Plan” offers a 15% discount for power usage between 7pm – 9am if most of your usage happens during that time. If not, then you lose your discount and you pay a 20% premium on the usage outside of that time. The cost per kilowatt hour is $0.05432. How much would you have paid in March if you used 450 KWH 7pm – 9am and 465 KWH 9am – 7pm?

OpenStudy (anonymous):

One of these is the answer- $49.70 $51.08 $54.75 $42.25

OpenStudy (anonymous):

have you done any work?

OpenStudy (anonymous):

hint, answer is between 50 and 52

OpenStudy (anonymous):

I tried, but I never done any of these so I had no idea.

OpenStudy (anonymous):

no worries so we will try to solve it, for the hours 7-9 we have CONSUMPTION=450 and cost for 1KWH=0.05432 so total cost for 7-9 is = 0.05432*450 but we get a discount too and that is 15%, so our cost will be 0.05432*450*0.85=20.77

OpenStudy (anonymous):

and for hours 9 to 7 we have a premium of 20% , cost = 1.2*465*0.05432=30.3

OpenStudy (anonymous):

add them to find total cost which should approx 51.2

OpenStudy (anonymous):

Ohhh! Okay, thank you!

OpenStudy (anonymous):

make sure you understand why I have use 1.2 and 0.85 for premium and discount

OpenStudy (anonymous):

15% off so it's 0.85 but idk about the 1.2.

OpenStudy (anonymous):

I see the 20% but Idk what the 1.2 is for quite yet..

OpenStudy (anonymous):

well we have 20% premium , right so say i buy a dvd for 100 dollars and it hs 20 percent premium, what is your finall cost you will say 20*100 /100=20= which is ON TOP OF 100 dollars(PREMIUM) SO total cost = 100+premium=120

OpenStudy (anonymous):

similarly for our question too and when it is a discount you subtract , so say i buy dvd with 20% discount my discount value = 20*100/100= 20=discount final cost = cost - discount = 100-20=80 or if you want to use a smart trick you can say final cost is 0.80*100=80

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

Ohh, alright, I see what you did there, thank you, that helps a lot.